HOT AND DIFFUSE CLOUDS NEAR THE GALACTIC CENTER PROBED BY METASTABLE H_{3}^{+}

\underline{TAKESHI OKA}, Department of Astronomy and Astrophysics and Department of Chemistry, University of Chicago; THOMAS R. GEBALLE, Gemini Observatory, Hilo, Hawaii; MIWA GOTO, Max Planck Institute for Astronomy, Heidelberg, Germany; TOMONORI USUDA, Subaru Telescope, National Astronomical Observatory of Japan, Hilo, Hawaii; BENJAMIN J. McCALL, Department of Chemistry and Department of Astronomy, University of Illinois at Urbana-Champaign.

We have observed a vast amount of high temperature (T ∼ 250 K) and low density (n ∼ 100 cm^{−3}) gas with a large velocity dispersion in the Central Molecular Zone (CMZ) of the Galactic center. We used H_{3}^{+} which is a sensitive probe of low density molecular gas. The observed large column density of H_{3}^{+} in the (3, 3) metastable rotational level gives evidence for high temperature, and absence in the (2, 2) level indicates low density. This remarkable non-thermal rotational distribution is caused by metastability of the (3, 3) level and the fast (2, 2) → (1, 1) spontaneous emission (27 days)^{a}. The strongest absorption component observed toward the bright infrared source GCS 3-2 is at velocity of ∼ - 100 km s^{−1}, indicating that about a half of the hot and diffuse gas is associated with the 180 pc Expanding Molecular Ring. The other half with lower velocities of - 50 km s^{−1} and ∼ 0 km s^{−1} is closer to the Galactic center. The large H_{3}^{+} column density indicates high ionization rate on the order of 10^{−14} s^{−1} in the CMZ if the C/H ratio is indeed as high as reported. With the hot X-rays and high magnetohydrodynamic activities, such a high value may be reasonable. The non-thermal rotational distribution of H_{3}^{+} has also been observed toward 7 other infrared sources within 30 pc of the Galactic center indicating that the hot and diffuse gas is ubiquitous in the CMZ. The spectrum toward GC IRS 3 near Sgr A* shows presence of the hot and diffuse gas in the “50 km s^{−1} cloud”, the complex of giant molecular clouds which plays a central role in the discussion of Sgr A* and its environment. Of the observed total H_{3}^{+} column density toward GCS 3-2 of 4.3 × 10^{15} cm^{−2}, approximately 3.1 × 10^{15} cm^{−2} is inferred to be in the CMZ while 1.2 × 10^{15} cm^{−2} is in the intervening spiral arms. Almost all of H_{3}^{+} in the CMZ is in diffuse clouds. This suggests that the previously reported volume filling factor (f ≥ 0.1) of dense clouds is an overestimate by at least an order of magnitude.

aT. Oka and E. Epp, ApJ, 613, 349 (2004)